Simply Chocolate Company is considering two possible expansion plans.Proposal X involves opening five stores in North Carolina at a cost of $2,400,000. Under Proposal Y, the company would focus on Virginia and open six stores at a cost of $3,000,000. The following information is given for the two proposals:

                                                                 

                                                                           Proposal X           Proposal Y

Required investment                                      $2,400,000        $3,000,000

Estimated life                                                   10 years              10 years

Estimated residual value                                  $200,000              $200,000

Estimated annual net cash flows                        $450,000             $580,000

Required rate of return                                       14%                      14%

Based on the above following problem,

Required: for each proposal, you are asked to calculate

a.Pay back Period

b.Accounting Rate of Return

c.Net Present Value      

d.Profitability Index

3.Indicate which proposal is the better investment.
Answer #1

a. Payback period for Proposal X= 2400000/450000=5.33 year and for proposal Y=3000000/580000= 5.17 year

b. Accounting rate of return for X=450000/2400000=18.75% and for Y= 580000/3000000=19.33%

c.

For Proposal X present value of annual net cash flow=$2347252.04.Calculation given below:

A C D E G rate 14% 14% rate PMT 450000 PMT 450000 10 10 nper nper PV |=PV(E1,E3,E2) ($2,347,252.04) PV PV(rate, nper, pmt, [f

So, NPV=-2400000+2347252.04+200000/1.14^10=$1200.80

For proposal Y present value of annual net cash flow=$3025347.07 Calculation given below:

A B D F rate PMT 14% 14% rate 580000 PMT 580000 nper PV 10 10 nper |=PV(E1,E3,E2) ($3,025,347.07) PV PV(rate, nper, pmt, [fv]

So, NPV for proposal Y= 3025347.07-3000000+200000/1.14^10=$79295.83

d. Profitability index for X= (NPV+Initial investment)/initial investment=(1200.80+2400000)/2400000=1.0005

Similarly PI for Y= (79295.83+3000000)/3000000= 1.026

As PI for proposal Y is more than X, so proposal Y should be chosen.